What is the y-intercept of the line containing (2,7) and perpendicular to a line with slope -3/2?
Answer by mananth(16946)

(2,7) and perpendicular to a line with slope -3/2?
slope of perpendicular line will be 2/3 ( negative reciprocal)
point (2,7)
m= 2/ 3

Plug value of the slope and point (2,7) in
Y = m x + b
7.00 = 4/3 + b
b= 7 - 4/3
b= 17/ 3
So the equation will be
Y = 2/3 x + 17/3
17/3 is the intercept

m.ananth@hotmail.ca